OpenStudy (anonymous):

what is the solution to x + (-10) = 38

OpenStudy (akashdeepdeb):

Add 10 to both sides! Like this.... x -10 + 10 = 38 + 10 Understood? :)

OpenStudy (uri):

Add 10 both sides x + (-10) = 38 x-10=38 10 10 ========= @kaylajasmine80 what will be left?